OpenStudy (anonymous):

1. Which property is used in the statement? If y = 3 and 3 = 2 + 1, then y = 2 + 1. transitive property symmetric property addition property of equality subtraction property of equality 2. Which property is used in the statement? If 2w = –8, then w = –4. reflexive property symmetric property division property of equality subtraction property of equality 3. Which pattern could have been used to create the sequence? 400, –200, 100, –50, 25

OpenStudy (anonymous):

I will fan anyone and give them a mdl if they help me.

OpenStudy (anonymous):

@Michele_Laino please help!

OpenStudy (michele_laino):

question #1: the property: "if a=b and b=c, then a=c" is the transitive property

OpenStudy (anonymous):

thank you so much what about the other ones?

OpenStudy (michele_laino):

Question #2: if we have the subsequent equation: \[5w = - 15\] I can divide both sides by 5, and I get: \[w = - 3\]

OpenStudy (anonymous):

ok so which one is it?

OpenStudy (michele_laino):

hint: I have "divided" both sides

OpenStudy (michele_laino):

Question #3: we have: -200= 400/(-2) 100=-200/(-2) -50=100/(-2) and so on....

OpenStudy (anonymous):

oh ok te division so for the 3thrd question the rest is Multiply the previous term by –2 Divide the previous term by –2 Subtract 600 from the previous term. Add –200 to the previous term

OpenStudy (anonymous):

which one is it?

OpenStudy (michele_laino):

please look at my previous post

OpenStudy (anonymous):

what is it I don't understand @Michele_Laino

OpenStudy (michele_laino):

here it is: Question #3: we have: -200= 400/(-2) 100=-200/(-2) -50=100/(-2) and so on....

OpenStudy (michele_laino):

every time I have divided the previous term by the same quantity

OpenStudy (anonymous):

so is it? Add –200 to the previous term

OpenStudy (michele_laino):

no, since I have divided by -2